# Simpler Hamiltonian structure of Gravity in degenerate limit

### v3

@@ Abstract
We analyze Hamiltonian gravity in the limit of a vanishing (spatial) triad determinant. In the context of Barbero-Immirzi formulation, the Lorentzian constraints in this limit are found to exhibit a formal similarity to the Euclidean constraints of Ashtekar’s self-dual formulation, modulo the difference that the spatial diffeomorphisms become trivial. Further, we show that the Carrollian limit of gravity emerges as a special case, thus providing it a geometric interpretation independent of either the speed of light or coupling constants.
@@ I Introduction
The obstacles to a canonical quantization of Lorentzian gravity are fairly well-known, irrespective of whether the canonical variables are defined within a metric or a connection formulation. The structure of the Hamiltonian constraint itself, containing both extrinsic and spatial curvature, poses a serious challenge given the inherent operator ordering ambiguity, non-polynomiality and regularization issues. The hunt for a simpler Hamiltonian structure have also led to efforts towards constructing different limits of gravity. It could be worthwhile to mention a few of these here, such as the strong coupling limit [1, 2], $G\rightarrow 0$ limit [3, 4], $G\rightarrow\infty$ limit [5], the topological limit [6] and so on.
@@
In the Hamiltonian theory of Hilbert-Palatini gravity based on the connection formalism, the densitized triad fields emerge as a natural canonical variable (momenta). In a quantum theory, a discrete quantum geometry could be anticipated, where trivial densitized triads could potentially define the lowest eigenstates provided the other (regularized) constraints are also satisfied. Hence, it is worthwhile to explore classical gravity first in a limit where the spatial geometry becomes trivial in the sense of densitized triads and the determinant of the four-metric. Such a limit for gravity has not been explicitly constructed so far.
@@
To this end, in this brief note we explore a limit where the determinant of the spatial triad (and also of the four-metric) vanishes. We implement the limit by defining a local scaling transformation on the canonical variables within the connection formalism.
@@
First, we apply this limit to the Barbero-Immirzi Hamiltonian formulation [7, 8]. We elucidate two different reductions of the constraints based on two different degenerate limits introduced here. In both cases, the resulting Hamiltonian theories reflect remarkable simplications. In the first case, we find that the Lorentzian constraints exhibit a formal similarity to the Euclidean constraints of Ashtekar’s self-dual gravity [9, 10, *ashtekar1], modulo differences regarding the action of spatial diffeomorphism. As an example, we find a solution to this limiting gravity theory in the context of inhomogeneous and anisotropic cosmology. In case of the second limit, we find a Hamiltonian structure whose quantum counterpart would be free of any operator ordering ambiguity. Next, we find that a special case of the degenerate limit corresponds to Carrollian gravity, the latter being originally defined for a vanishingly small speed of light [12, 13, 1, 2, 14, 15].
Apart from exploring possible new limits of Hamiltonian gravity, our analysis could provide a fresh perspective towards a perturbative expansion of gravity around vanishing densitized triad, different from the standard approach of expanding around a fixed background geometry with $G$ being the perturbative parameter.
@@
In the next section, we first set up the relevant notation through a very brief outline of Barbero-Immirzi canonical gravity. The limit to a degenerate tetrad through a set of scaling transformations is then introduced. The implications of this limit for the Hamiltonian theory is discussed, along with an explicit example. In Sec-III we explore a different limit associated with a degenerate tetrad. This is followed by an analysis of the special case of global scaling transformations, the resulting theories being shown to be equivalent to the Carrolian gravity theories. We conclude our analysis with a few relevant remarks. In the Appendix, as a relevant digression, we present a corollary on Birkhoff’s theorem in ‘electric’ Carroll gravity based on the associated Hamiltonian structure.
@@ II Degenerate limit of Barbero-Immirzi gravity
${\cal L}(e,\omega)=\frac{1}{2\kappa}\Big{[}ee^{\mu}_{I}e^{\nu}_{J}R_{\mu\nu}^{~{}~{}~{}IJ}(\omega)~{}+~{}\eta\epsilon^{\mu\nu\alpha\beta}\Big{(}D_{\mu}e_{\nu}^{I}D_{\alpha}e_{\beta I}-e_{\mu}^{I}e_{\nu}^{J}R_{\alpha\beta}^{~{}~{}~{}KL}(\omega)\Big{)}\Big{]}$ (1)
Assuming the standard parametrization of the tetrad as presented in Ref.[17, 18], the totally constrained Hamiltonian density reads:
${\cal H}=NH+N^{a}H_{a}+\frac{1}{2}\omega_{t}^{~{}IJ}G_{IJ}.$ (2)
The scalar Lapse $N$, the shift $N^{a}$ and the temporal connection fields $\omega_{t}^{~{}IJ}$ emerge as Lagrange multipliers above, whereas $H\approx 0,~{}H_{a}\approx 0,~{}G_{IJ}\equiv(G_{0i},~{}G_{ij})\approx 0$ denote the Hamiltonian, spatial diffeomorphism, boost and rotation constraints, respectively. Their explicit expressions are presented later. The spatial metric is defined as: $q_{ab}~{}:=~{}e_{a}^{I}e_{bI}$, and the tetrad determinant as: $e:=det(e^{I}_{\mu})=N\sqrt{q}$.
@@
where the conjugate pair is given by: $A_{a}^{i}:=\omega_{a}^{0i}+\frac{1}{2}\epsilon^{ijk}\omega_{a}^{jk},~{}E^{a}_{i}:=\pi^{a}_{~{}0i}=ee^{a}_{i},~{}q=\det E^{a}_{i}:=E$. The connection fields $A_{a}^{i}$ emerge as the $SU(2)$ gauge field. In terms of this canonical pair, the set of constraints read [19, 16]:
$G^{rot}_{i}~{}$ $=$ $~{}-\eta D_{a}(A)E^{a}_{i},$
$H_{a}~{}$ $=$ $~{}E^{b}_{i}F_{ab}^{i}(A),$
$H~{}$ $=$ $~{}\frac{\kappa}{2\sqrt{E}}E^{a}_{i}E^{b}_{j}\Big{[}\eta\epsilon^{ijk}F_{ab}^{k}(A)-(1+\eta^{2})\Big{(}\partial_{[a}\omega_{b]}^{~{}ij}(E)+\omega_{[a}^{~{}ik}(E)\omega_{b]}^{~{}kj}(E)\Big{)}\Big{]},$ (4)
@@
Limit to degenerate tetrad
@@
The brackets in the spatial label ‘$(a)$’, which can run over the three space coordinates, imply that it is not summed over when repeated. The primed variables are assumed to remain finite in the limit. To emphasize, each of the three vectors for a fixed $(a)$ need not vanish in this limit. The triad determinant $\sqrt{q}=\delta\sqrt{q^{\prime}}$ goes to zero as $\delta=\prod_{a}\delta_{(a)}\rightarrow 0$, which defines the ‘degenerate’ limit. This in turn implies that: $\delta_{(a)}\delta_{(b)}=\frac{\delta}{\delta_{(c)}}\rightarrow 0$ for any $a\neq b$ in the limit. An explicit realization of these properties is presented through an example later.
@@
$G^{rot}_{i}$ $=$ $\epsilon^{ijk}A_{a}^{{}^{\prime}k}E^{{}^{\prime}a}_{j}-\eta\frac{\delta}{\delta_{(a)}}\partial_{a}E^{{}^{\prime}a}_{i}-\eta\partial_{a}\frac{\delta}{\delta_{(a)}};$
@@
-- $E^{{}^{\prime}b}_{i}A_{b}^{{}^{\prime}i}\frac{\delta}{\delta_{(a)}}\frac{\delta_{(b)}}{\delta}\partial_{a}\frac{\delta}{\delta_{(b)}}+E^{{}^{\prime}b}_{i}A_{a}^{{}^{\prime}i}\frac{\delta}{\delta_{(b)}}\frac{\delta_{(a)}}{\delta}\partial_{a}\frac{\delta}{\delta_{(a)}}\Big{]},$
$H$ $=$ $\frac{\kappa}{2\delta\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}A_{a}^{{}^{\prime}i}A_{b}^{{}^{\prime}j}~{}+~{}\delta^{2}\Big{(}\frac{1}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}+\omega_{a}^{{}^{\prime}ik}\omega_{b}^{{}^{\prime}kj}\Big{)}+\omega^{{}^{\prime}ij}_{b}\frac{\delta^{2}}{\delta_{(a)}\delta_{(b)}}\partial_{a}\frac{\delta_{(b)}}{\delta}\Big{]}$
@@
$H^{\prime}_{a}$ $=$ $\frac{1}{\eta}A_{a}^{{}^{\prime}}G^{{}^{\prime}rot}_{i};$
@@
Noting that the $SU(2)$ covariant derivative and the field-strength in the limit are given by: $D^{\prime}_{a}(A^{\prime})E^{{}^{\prime}a}_{i}=-\frac{1}{\eta}\epsilon^{ijk}E^{{}^{\prime}b}_{j}A_{a}^{{}^{\prime}j}E^{{}^{\prime}ak},~{}F_{ab}^{{}^{\prime}~{}i}=-\frac{1}{\eta}\epsilon^{ijk}A^{{}^{\prime}j}_{a}A_{b}^{{}^{\prime}k}$, the above set may be rewritten as:
$G^{{}^{\prime}rot}_{i}$ $=$ $\eta D^{\prime}_{a}(A^{\prime})E^{{}^{\prime}a}_{j};$
@@
This form precisely resembles the set of Euclidean constraints of Ashtekar’s self-dual gravity (which could be obtained by setting $\eta=i$ in Barbero-Immirzi formulation). The important difference, however, is that the Lorentzian set here corresponds to a simpler canonical structure, given that the spatial diffeomorphism constraints are trivially satisfied.
@@
where we assume that: $\alpha+\beta>0,~{}\beta+\gamma>0,~{}\alpha+\gamma>0$. A similar metric was first considered by BKL [20, *bkl1], which could be a solution of Einstein gravity in presence of matter. Assuming that the effect of matter could be neglected in the degenerate limit (which here represents a curvature singularity at $t=0$), we proceed to find the associated solution.
Let us first demonstrate that the assumed generic properties of the scaling transformation are all satisfied. The metric degenerates as $t\rightarrow 0$. This limit could be parametrized as $\epsilon\rightarrow 0$ where $t=\epsilon t^{\prime}$ ($t^{\prime}$ being finite). Inserting this in the metric solution just obtained above, we find:
@@
As the above solution reflects, the spatial dependence of these exponents are arbitrary. Thus, the spatial points are uncorrelated and at each point there could be an independent set of Kasner exponents. In other words, the degenerate limit in this context leads to a classical realization of ultralocal physics.
III A different degenerate limit: Hamiltonian structure without ordering ambiguity

### v4

# Canonical Gravity in Degenerate Limit

@@ Abstract
We construct a limit of Hamiltonian gravity as the determinant of the spatial triad (and hence of the four-metric) goes to zero. Within the Barbero-Immirzi $SU(2)$ formulation, we present two possible realizations of this limit, with the consequence that the Hamiltonian constraint becomes simpler and spatial diffeomorphisms become trivial. In the first case, the Hamiltonian constraint exhibits a polynomial structure, being formally similar to the Euclidean Hamitonian constraint of Sen-Ashtekar self-dual formulation. In the latter, the constraints become free from ordering ambiguity. Further, we show that the Carrollian gravity emerges as a special case of this degenerate limit, thus providing it a new geometric interpretation independent of the speed of light or any dimensionful coupling constant ($G$).
@@ I Introduction
The obstacles to a canonical quantization of Lorentzian gravity are well-known, irrespective of whether the canonical variables are defined within a metric or a connection formulation [1, 2]. The structure of the Hamiltonian constraint itself, containing both extrinsic and spatial curvature, poses a serious challenge given the inherent operator ordering ambiguity owing to the appearance of non-commuting operators, non-polynomiality in the canonical variables and regularization issues due to presence of coincident operators or distributions [3, 4, 5, 6]. The hunt for a simpler Hamiltonian structure have also led to efforts towards formulating different limits of gravity, with an aim to gain insights exploiting the potential simplifications. This endeavour is reflected in constructions such as the strong coupling limit [7, 8, 9, 10], $G\rightarrow\infty$ limit [11], $G\rightarrow 0$ limit [12, 13], the topological limit [14] and so on.
@@
In the Hamiltonian theory of Hilbert-Palatini gravity based on the connection formalism, the densitized triad fields emerge as a natural canonical variable (momenta) [1, 2, 5, 15]. In a quantum theory, a discrete quantum geometry could be anticipated, where trivial densitized triads could potentially define the lowest eigenstates provided the other (regularized) constraints are also satisfied. Hence, it is worthwhile to explore classical gravity first in a limit where the spatial geometry becomes trivial in the sense of densitized triads having a vanishing determinant. Under certain reasonable assumptions on the lapse function, this implies that the determinant of the four-metric should also vanish in such a limit. Even though discussions on degenerate extensions of general relativity have appeared in the earlier literature [16, 17, *romano, 19, *sengupta2016, 21], such a limit for gravity has not been explicitly constructed so far.
Analysis of gravity theory in a limit of degenerate metric (tetrad) is also important in order to understand the singularities better. In particular, such a construction could be useful whenever an approach to a singularity is directly associated with a degenerating metric. For instance, such a scenario is realized for spacelike singularities (e.g. cosmological singularity [22, 23]) where spatial metric determinant approaches a zero. To emphasize, the dynamics resulting from the equations of motion in this limit need not be equivalent to the dynamics far away from singularity. This could provide new insights as to whether gravity behaves differently close to a singularity which is approached through a limit to degenerate metrics.
@@
To this end, in this note we explore a limit where the determinant of the spatial triad (and also of the four-metric) vanishes. We implement the limit by defining a local scaling transformation on the canonical variables within the connection formalism.
@@
First, we apply this limit to the Barbero-Immirzi formulation of gravity [24, 25]. We obtain two different reductions of the associated constraints. In both cases, the resulting Hamiltonian theories reflect remarkable simplications. In the first case, we find that the Lorentzian constraints exhibit a formal similarity to the Euclidean constraints of Sen-Ashtekar self-dual gravity [26, 27, *ashtekar1], modulo differences regarding the action of spatial diffeomorphism. As an example, we find a solution to this limiting gravity theory in the context of inhomogeneous and anisotropic cosmology. In case of the second limit, we find a Hamiltonian structure whose quantum counterpart would be free of any ordering ambiguity associated with non-commuting operators. Next, we find that a special case of the degenerate limit corresponds to Carrollian gravity [7, 8, 29, 30], the latter being originally defined for a vanishingly small speed of light [31, 32].
Apart from exploring possible new limits of Hamiltonian gravity, our analysis could provide a fresh perspective towards the issue of solving the spatial diffeomorphism and the Hamiltonian constraint. To emphasize, we show that the degenerate limit(s) provide a way to solve the spatial diffeomorphism classically, before proceeding towards a canonical quantization. Along with the simpler structure of the Hamiltonian constraint reflected in its polynomial form in particular, achieved without taking recluse to Euclidean gravity, this feature seems promising from the perspective of canonical quantization.
@@
In the next section, we first set up the relevant notation through a very brief outline of Barbero-Immirzi canonical gravity. The limit to a degenerate tetrad through a set of scaling transformations is then introduced. The implications of this limit for the Hamiltonian theory is discussed, along with an explicit example. In Sec-III we explore a different limit associated with a degenerate tetrad. This is followed by an analysis of the special case of global scaling transformations, leading to theories equivalent to Carrollian gravity. We conclude our analysis with a few relevant remarks. Some basic details regarding the symplectic form of Barbero-Immirzi real SU(2) formulation is provided in Appendix A. In Appendix B, as a relevant digression, we present a corollary on Birkhoff’s theorem in ‘electric’ Carroll gravity based on the associated Hamiltonian structure.
@@ II Degenerate limit of Barbero-Immirzi gravity
${\cal L}(e,\omega)=\frac{1}{2\kappa}\Big{[}ee^{\mu}_{I}e^{\nu}_{J}R_{\mu\nu}^{~~~IJ}(\omega)~+~\eta\epsilon^{\mu\nu\alpha\beta}\Big{(}D_{\mu}e_{\nu}^{I}D_{\alpha}e_{\beta I}-e_{\mu}^{I}e_{\nu}^{J}R_{\alpha\beta IJ}(\omega)\Big{)}\Big{]}$ (1)
The $SO(3,1)$ field strength above is defined as: $R_{\mu\nu}^{~~~IJ}(\omega)=\partial_{[\mu}\omega_{\nu]}^{~IJ}+\omega_{[\mu}^{~IK}\omega_{\nu]K}^{~~~J}$. Assuming the standard parametrization of the tetrad as presented in Ref.[34, 15], the totally constrained Hamiltonian density reads:
${\cal H}=\frac{1}{2}\omega_{t}^{~IJ}G_{IJ}+N^{a}H_{a}+NH.$ (2)
The scalar Lapse $N$, the shift $N^{a}$ and the temporal connection fields $\omega_{t}^{~IJ}$ emerge as Lagrange multipliers above, whereas $H\approx 0,~H_{a}\approx 0,~G_{IJ}\equiv(G_{0i},~G_{ij})\approx 0$ denote the Hamiltonian, spatial diffeomorphism, boost and rotation constraints, respectively. Their explicit expressions are presented later. The spatial metric is defined as: $q_{ab}~:=~e_{a}^{I}e_{bI}$, and the tetrad determinant as: $e:=det(e^{I}_{\mu})=N\sqrt{q}$. Further relevant details in this context are provided in Appendix A.
@@
where the conjugate pair is given by: $A_{a}^{i}:=\omega_{a}^{0i}+\frac{\eta}{2}\epsilon^{ijk}\omega_{a}^{jk},~E^{a}_{i}:=\pi^{a}_{~0i}=ee^{a}_{i},~q=\det E^{a}_{i}:=E$. The connection fields $A_{a}^{i}$ emerge as the $SU(2)$ gauge field. The momenta $E^{a}_{i}$ are known as the densitized spatial triad fields. These conjugate fields are analogous to the gauge potential and the electric fields of $SU(2)$ Yang-Mills theory, respectively.
In terms of this canonical pair in the symplectic form, the set of constraints read [35, 33]:
$G^{rot}_{i}$~ $=$ $~-\eta D_{a}(A)E^{a}_{i}\approx 0,$
$H_{a}$~ $=$ $~E^{b}_{i}F_{ab}^{i}(A)+\Big{(}\eta+\frac{1}{\eta}\Big{)}\Bigg{(}A_{a}^{i}+\frac{\eta}{2}\epsilon^{ijk}\omega_{a}^{~jk}(E)\Bigg{)}G^{rot}_{i}\approx 0,$
$H$~ $=$ $~\frac{\kappa}{2\sqrt{E}}E^{a}_{i}E^{b}_{j}\Big{[}\eta\epsilon^{ijk}F_{ab}^{k}(A)-(1+\eta^{2})\Big{(}\partial_{[a}\omega_{b]}^{~ij}(E)+\omega_{[a}^{~ik}(E)\omega_{b]}^{~kj}(E)\Big{)}\Big{]}\approx 0,$ (4)
@@
These three constraints above are the generators of the symmetry transformations associated with the spatial rotation, spatial diffeomorphisms (modulo rotations) and temporal translations, respectively.
A limit to degenerate tetrad
@@
The brackets in the spatial label ‘$(a)$’, which can run over the three space coordinates, imply that it is not summed over when repeated. The primed variables are assumed to remain finite in the limit. To emphasize, each of the three vectors for a fixed $(a)$ need not vanish in this limit. The triad determinant $\sqrt{q}=\delta\sqrt{q^{\prime}}$ goes to zero as $\delta=\prod_{a}\delta_{(a)}\rightarrow 0$, which defines the ‘degenerate’ limit. This in turn implies that: $\delta_{(a)}\delta_{(b)}=\frac{\delta}{\delta_{(c)}}\rightarrow 0$ for any $a\neq b\neq c$ in the limit. An explicit realization of these properties is presented through an example later.
@@
$G^{rot}_{i}$ $=$ $\epsilon^{ijk}A_{a}^{{}^{\prime}k}E^{{}^{\prime}a}_{j}-\eta\frac{\delta}{\delta_{(a)}}\partial_{a}E^{{}^{\prime}a}_{i}-\eta E^{{}^{\prime}a}_{i}\partial_{a}\frac{\delta}{\delta_{(a)}};$
@@
- $E^{{}^{\prime}b}_{i}A_{b}^{{}^{\prime}i}\frac{\delta}{\delta_{(a)}}\frac{\delta_{(b)}}{\delta}\partial_{a}\frac{\delta}{\delta_{(b)}}+E^{{}^{\prime}b}_{i}A_{b}^{{}^{\prime}i}\frac{\delta}{\delta_{(b)}}\frac{\delta_{(a)}}{\delta}\partial_{a}\frac{\delta}{\delta_{(a)}}+\Big{(}\eta+\frac{1}{\eta}\Big{)}\Bigg{(}A_{a}^{{}^{\prime}i}+\frac{\eta}{2}\epsilon^{ijk}\omega_{a}^{{}^{\prime}jk}\Bigg{)}G^{{}^{\prime}rot}_{i}\Big{]}$
$H$ $=$ $\frac{\kappa}{2\delta\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}A_{a}^{{}^{\prime}i}A_{b}^{{}^{\prime}j}~+~\delta^{2}\Big{(}\frac{1}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}+\omega_{a}^{{}^{\prime}ik}\omega_{b}^{{}^{\prime}kj}\Big{)}+\omega^{{}^{\prime}ij}_{b}\frac{\delta^{2}}{\delta_{(a)}\delta_{(b)}}\partial_{a}\frac{\delta_{(b)}}{\delta}\Big{]}$ (8)
@@
$H^{\prime}_{a}$ $=$ $\eta A_{a}^{{}^{\prime}i}G^{{}^{\prime}rot}_{i};$
@@
Noting that the $SU(2)$ covariant derivative and the field-strength in the limit are given by: $D^{\prime}_{a}(A^{\prime})E^{{}^{\prime}a}_{i}=-\frac{1}{\eta}\epsilon^{ijk}A_{a}^{{}^{\prime}j}E^{{}^{\prime}ak},~F_{ab}^{{}^{\prime}~i}=-\frac{1}{\eta}\epsilon^{ijk}A^{{}^{\prime}j}_{a}A_{b}^{{}^{\prime}k}$, the above set may be rewritten as:
$G^{{}^{\prime}rot}_{i}$ $=$ $\eta D^{\prime}_{a}(A^{\prime})E^{{}^{\prime}a}_{i};$
@@
Note that modulo an overall density factor which could be absorbed away through a densitized lapse function, the Hamiltonian constraint has now become a quadratic polynomial in the canonical variables.
Let us observe that this same feature arises in the Sen-Ashtekar self-dual connection formulation, which, however, corresponds to Euclidean gravity (and could be obtained by setting $\eta=i$ in Barbero-Immirzi formulation). The other important difference is that the Lorentzian set here corresponds to a simpler (and weaker) canonical structure, given that the spatial diffeomorphism constraints are trivially satisfied.
It is also straightforward to observe that the reduced constraints preserve the $SU(2)$ interpretation. For instance, the rotation constraint generates the following transformation of the canonical pair:
$\delta A_{a}^{i}=[A_{a}^{i},\int\theta^{k}G_{k}^{rot}]=\epsilon^{ijk}A_{a}^{j}\theta^{k}=\eta D_{a}\theta^{i},$
$\delta E^{a}_{i}=[E^{a}_{i},\int\theta^{k}G_{k}^{rot}]=\epsilon^{ijk}\theta^{k}E^{a}_{j}$ (11)
using the fact that in the limit the gauge-covariant derivative reduces to: $D_{a}\theta^{i}=\frac{1}{\eta}\epsilon^{ijk}A_{a}^{j}\theta^{k}$. These are manifestly same as the standard $SU(2)$ gauge transformations acting on the gauge potential and the electric field components, respectively.
@@
where we assume that: $\alpha+\beta>0,~\beta+\gamma>0,~\alpha+\gamma>0$. A similar metric was first considered by BKL [22, *bkl1], which could be a solution of Einstein gravity in presence of matter. Assuming that the effect of matter could be neglected in the degenerate limit (which here represents a curvature singularity at $t=0$), we proceed to find the associated solution.
Let us first demonstrate that the assumed generic properties of the scaling transformation are all satisfied. The spatial metric (as well as the four-metric) degenerates as $t\rightarrow 0$. This limit could be parametrized as $\epsilon\rightarrow 0$ where $t=\epsilon t^{\prime}$ ($t^{\prime}$ being finite). Inserting this in the metric solution just obtained above, we find the components of $\delta_{(a)}$ defined in eq.(6) as:
@@
As the above solution reflects, the spatial dependence of these exponents are arbitrary, unlike in the case of Kasner solution of vacuum Einstein gravity where these are forced to be constants by the field equations. Thus, the spatial points are uncorrelated and at each point there could be an independent set of Kasner exponents. In other words, the degenerate limit in this context leads to a classical realization of ultralocal physics.
III An alternative limit: Hamiltonian structure without ordering ambiguity